OpenStudy (anonymous):

Graph the line passing through the given point and having the given slope. Give the slope-intercept form of the equation of the line if possible. (1,-2),m=-3/5 Can anyone help???

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

slope-intercept form of the line: y-y1=m(x-x1) you have all data that you need..

OpenStudy (anonymous):

y-(-2) = -3/5 *(x-1) simplify :)

OpenStudy (anonymous):

Thank you very much for your answer but I need to see the step by step instructions because I do not understand how to do these type problems. Would anyone be willing to share step by step instructions?